# GNN-Transformer Cooperative Architecture for Trustworthy Graph Contrastive Learning

### v1

@@ Conclusion
Most of the existing GCL methods utilize graph augmentation strategies, which may perturb the underlying semantics of graphs. Furthermore, they utilize GNNs as graph encoders, which inevitably results in the occurrence of over-smoothing and over-squashing issues. To address these issues, we propose GNN-Transformer Cooperative Architecture for Trustworthy Graph Contrastive Learning (GTCA). GTCA uses GNNs and NodeFormer as encoders to generate node representation views. In addition, it utilizes topological property of graphs to generate the topology structure views. Theoretical analysis and experimental results demonstrate the effectiveness of GTCA.

### v2

@@ Conclusion
Most of the existing GCL methods utilize graph augmentation strategies, which may perturb the underlying semantics of graphs. Furthermore, they utilize GNNs as graph encoders, which inevitably results in the occurrence of over-smoothing and over-squashing issues. To address these issues, we propose GNN-Transformer Cooperative Architecture for Trustworthy Graph Contrastive Learning (GTCA). GTCA uses GCN and NodeFormer as encoders to generate node representation views. In addition, it utilizes topological property of graphs to generate the topology structure views. Theoretical analysis and experimental results demonstrate the effectiveness of GTCA.